OpenStudy (itsmichelle29):

The function t(x) = x + 6 determines how many cans of soup a food truck needs to stock, where x is the number of shifts the crew is going to work in the truck. The crew uses c(t(x)) to find the amount of money to spend on soup. The function c(x) = 2x + 4. Solve for how much money must be spent when the crew is going to work 4 shifts. 24 28 30 34

OpenStudy (itsmichelle29):

is it 24???

OpenStudy (lynfran):

t(x)=x+6 and 4 shift...t(4)=4+6=10... c[t(x)]=2(10)+4

OpenStudy (itsmichelle29):

so it is 24??

OpenStudy (lynfran):

your correct

OpenStudy (itsmichelle29):

thanks can you help me with one more

OpenStudy (itsmichelle29):

What is the equation of the graph below? y = - (x - 3)^2 - 2 y = - (x + 2)^2 - 3 y = (x - 3)^2 - 2 y = (x + 2)^2 - 3

OpenStudy (itsmichelle29):

N.M i got the answer